OpenStudy (anonymous):

How do i take the antiderivative of (sinxcosx)^2?

OpenStudy (anonymous):

need any direction.. Maybe i missed this but im completely lost

OpenStudy (kainui):

I went to Wolfram and clicked Show Steps. http://www.wolframalpha.com/input/?i=integral%20(sinxcosx)%5E2%20dx&t=crmtb01 I hope that can help, I'm not really sure how to give you a good answer on this one, since it looks straight forward but apparently it isn't.

OpenStudy (myname):

First (sinx*cosx)^2 = sin^2 x*cos^2 x Then you can use the half angle formula for both sin^2 x and cos^2 x [sin^2 x= (1-cos2x)/2 && cos^2 x=(1+cos2x)/2] So, sin^2 x * cos^2 x dx = (1-cos2x)/2 *(1+cos2x)/2 --->[(a+b)(a-b)= a^2-b^2] '' '''''''''' =(1-cos^2 2x)/4 Now you can break this apart as (1/4 - cos^2 2x/4) dx The antiderivative of 1/4 is just x/4 And next we find the antiderivative of the cos^2 2x / 4 by again using the half angle identity that we used before. So this would be cos^2 2x / 4 = (1+cos 2 2x)/(2*4) =1+cos4x/8 Next, we have to find the antiderivative of (1+cos4x)/8 (1+cos4x)/8 dx = x/8+sin4x/32 ---> This is just by using the general antiderivative Therefore, the final answer would be =>x/4 - x/8 - sin4x/32 + C I hope this helps you.